Which of the following are possible roots of the function. Select all that apply

5x3 + 19x2- 29x + 5


a. +- 1/5

b. +- 19/29

c. +- 1

d. +- 5

e. +- 0

f. +- 25

g. +- 25/19

h. +- 29/19


1
Expert's answer
2021-05-02T15:59:50-0400

Using try and error method

Try (x-1)

x-1 = 0

x= 1

= 5(-1)³ +19(-1)² -29(-1) +5

= 0

Therefore, x-1 is a factor.

To get the other two factors, divide 5x³ +19x² -29x +5 by x-1 using long division.

This will give you 5x² +24x-5

Factorize 5x² +24x -5 =(5x-1)(x+5)

Equate these three factors to zero to give the roots of the function 5x³ +19x² -29x +5

x-1 = 0; x = 1

x+5 = 0; x = -5

5x-1 =0; x = 1/5